OpenStudy (anonymous):

f(x) = 3x + 9, g(x) = 5x2 how do i find (f + g)(x)

OpenStudy (anonymous):

Add the 2 functions together.

OpenStudy (whpalmer4):

\[(f+g)(x) = f(x)+g(x) = \]

OpenStudy (cggurumanjunath):